OpenStudy (anonymous):

H6P2 Phase inverter question

OpenStudy (anonymous):

VIN=Dv+VT VDD=(2*VIN)-VT gm1=K*(VIN-VT) also equal to gm2 Vth = vout = -(K*(VIN-VT)*vin)/(K*(VIN-VT)) Rth = 1/(K*(VIN-VT))